OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (-2, 6) and has a slope of -1? A. x + y = -4 B. x + y = 4 C. x - y = -4 D. x - y = 4

jimthompson5910 (jim_thompson5910):

Were you able to find the equation of the line in slope intercept form?

OpenStudy (anonymous):

No I am really bad at this type of math. I have trouble understanding.

jimthompson5910 (jim_thompson5910):

y = mx+b is the general equation of any line in slope intercept form

jimthompson5910 (jim_thompson5910):

we're given a slope of -1, so m = -1

OpenStudy (anonymous):

oh okay yeah i knew that.

jimthompson5910 (jim_thompson5910):

we want this line to go through (-2, 6) so x = -2 and y = 6

jimthompson5910 (jim_thompson5910):

y = mx+b y = -1*x+b 6 = -1*(-2)+b 6 = 2 + b 6 - 2 = b 4 = b b = 4 So we know m = -1 and b = 4 giving us the equation y = -x + 4 making sense so far?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

now we just need to move that x term over to the left side to get it into standard form so we just add x to both sides to do that y = -x + 4 y+x = -x + 4 + x x+y = 4

jimthompson5910 (jim_thompson5910):

As a check, we can plug (x,y) = (-2,6) into x+y = 4 x+y = 4 -2+6 = 4 4 = 4 ... works so the answer is confirmed

OpenStudy (anonymous):

i learned that last year but i was never refreshed this helped me a lot thank you so much!

jimthompson5910 (jim_thompson5910):

you're welcome